### v1

# VR²: A Co-Located Dual-Headset Platform for Touch-Enabled Human–Robot Interaction Research

Abstract.
Touch-rich human–robot interaction (HRI) is difficult to study: building and programming physical robots is costly and slow, while VR-based robot prototypes often remove physical contact or break the tight coupling between an agent’s body and the user’s felt touch. We present VR2VR, a co-located dual–VR-headset platform for HRI research in which a participant and a hidden operator share the same physical space while experiencing different virtual embodiments. The participant sees an expressive virtual robot that interacts face-to-face in a shared virtual environment. In real time, the robot’s upper-body gestures, head and gaze behaviors, and facial expressions are mapped from the operator’s tracked motion and face signals. Because the operator is physically co-present and calibrated into the same coordinate frame, the operator can also physically touch the participant, enabling the participant to perceive robot touch aligned with the robot’s hands; finger and hand motion are mapped to the robot using inverse kinematics to support precise contact. Beyond faithful motion retargeting for limb teleoperation, our VR2VR system supports experimental control by retargeting or selectively enabling nonverbal channels (e.g., head only vs. head+eyes vs. head+eyes+facial expressions) while keeping physical interaction constant. We detail the system design, calibration workflow, and safety considerations, and demonstrate the platform through a touch-based Wizard-of-Oz HRI study, illustrating how VR2VR lowers barriers for rapidly prototyping and rigorously evaluating embodied, touch-centric robot behaviors.
1. Introduction
While robots have made huge progress in producing smooth, lifelike motions for physical tasks, they still struggle to display appropriate behavior during human-robot interaction. This is particularly evident when robots need to share a space with or act in tight contact with humans, such as in the context of physical HRI (Farajtabar and Charbonneau, 2024) or service and assistive robotics. In such cases, even if the robot’s main purpose is physical support, this should be accompanied by a transparent and legible social behavior to increase acceptance and user experience (Nguyen et al., 2018; Seifi et al., 2024).
Social robotics has long investigated which cues and in which combinations play a role in perceiving a robot as a sociable companion. Nevertheless, how to realize the multimodal, flexible, and context-appropriate social behavior that makes physical robots acceptable and trustworthy, in our homes or – even closer – in our peripersonal space, remains an open question. Ideally, the robot’s social and functional behavior should be co-designed along with its hardware and iteratively assessed with users, but this is often possible only once a full hardware prototype and behavior implementation is available. On the other hand, even if focusing just on designing an interaction for a specific use case, comparing it to another embodiment and showing the influence of a specific appearance
With such VR platform enriched with social retargeting, behavior design and its compatibility with a specific robot appearance or hardware configuration can be investigated even before the robotic platform is ready.
Motion retargeting in robotics has been used to facilitate imitation learning or teleoperation and to transfer goal-driven human-like motion to robots, yet a social counterpart to transfer dynamic non-verbal cues and their combinations in context has been missing.
Here, we propose multimodal social (cue) retargeting. This method relies on a VR-VR setup to display, capture, and validate social behavior during a realistic interaction, allowing for safe and user-centric pHRI already at the design stage.
- In this framework, prospective users can have a more realistic experience of the interaction than with images/videos/vignettes, as often applied in HRI research
- Multimodal data from dyadic interactions can be collected in synch and in anonymized form and used for various purposes: data from the wizard can be used for training a system with appropriate behaviour, in terms of timed gestures, expressions, contacts, w.r.t. the current situation (e.g., serious/happy/empathic/…), data from the user can be used to evaluate behavioral correlated of user experience (e.g., hand motion data or face expressions/face action units as proxies of stress, distrust, discomfort)
- Data can be collected anywhere, VR headsets are easily transportable
2. Dual VR co-location system
VR2 is a co-located dual-headset platform for touch-enabled HRI studies (Figure 1). A participant and a hidden operator share the same physical space but see different views: the participant is immersed in VR with an expressive robot embodiment, while the operator uses passthrough MR to teleoperate the robot (Wizard-of-Oz) and can physically touch the participant to provide aligned haptic contact.
2.1. System features
2.1.1. Co-location and asymmetric views
Co-location via spatial anchor. Two VR devices share a common coordinate system via a Meta spatial anchor (panel 4 in Figure 1). This avoids marker-based calibration (e.g., ArUco) and supports quick setup in different rooms.
Runtime placement adjustment. During setup, the operator can further adjust the robot’s virtual placement (e.g., small translation/yaw offsets) through a GUI to match the physical arrangement (e.g., across a table).
Asymmetric rendering for control and immersion. The operator sees the real environment with the robot overlaid through passthrough MR, enabling continuous awareness of the participant’s posture and proximity while maintaining alignment for safe touch (panel 2 in Figure 1). The participant experiences a fully virtual environment and sees the robot and their own tracked virtual hands (panel 3 in Figure 1).
2.1.2. Communication, logging, and touch aids
Realtime streaming. Mapped robot control data from the operator headset—including arm/hand motion, head orientation, and face-related control parameters (Section 2.2)—are streamed to the participant headset via a WebSocket server.
Synchronized logging. In parallel, the server logs multimodal data for analysis: robot pose/control streams, both operator and participant facial expressions and hand joints, and participant-focused interaction events such as gazed robot body parts (e.g., head/hands/torso) and hand collisions with robot body parts.
Pre-recorded speech triggering. To support controlled verbal behavior without occupying the operator’s hands, VR2 integrates a foot pedal to trigger pre-recorded robot speech (panel 5 in Figure 1). The pedal provides three buttons and supports single- and double-press bindings, enabling quick navigation (e.g., previous/next clip) and branching between alternative speech-flow paths.
Touch alignment aids. Because the operator is co-located and calibrated into the shared frame, the operator can physically touch the participant while the participant sees corresponding robot touch. To better match the robot fingertip geometry during contact, the operator can wear 3D-printed fingertip covers shaped like Allegro fingertips (panel 6 in Figure 1).
2.2. Mapping human data to the robot
VR2 maps the operator’s tracked head, hands, and face signals to the robot embodiment in real time (Figure 2). The mapping has two goals: (1) enable expressive and legible nonverbal behaviors (gaze, facial affect proxies, head orientation), and (2) support touch alignment by making the robot’s visible hand/fingertip contacts coincide with the operator’s physical contacts with the participant.
2.2.1. Retargeting pipeline
Face and gaze screen-face controls. Quest Pro face tracking outputs 70 per-muscle blendshape values in .111https://developers.meta.com/horizon/documentation/unity/move-face-tracking We select seven channels that most affect perceived emotion in our screen-face embodiment: eye closedness , lip-corner dimple , brow lower (left/right), and chin raise ; the headset also provides gaze direction . Our robot face uses two stylized “eye” cylinders rendered on a screen; their shape and pose are controlled by parameters including eyelid-profile vertices and , eye rotation , eye depth scale , ear rotation (one DoF per ear), and on-screen eye position (Figure 2, step 1). We use a simple monotonic mapping that (i) closes/compresses the eyes with , (ii) modulates eye/ear “attitude” with combined brow/chin activity, and (iii) drives on-screen eye position from gaze with normalization and clamping; full equations are provided in Appendix A.
Head orientation retargeting. The headset provides a 6-DoF head pose. Because the robot base is fixed in the scene, we retarget only the operator’s head orientation to the robot head while discarding translation to keep the robot rooted.
Hands robot arms via IK for touch alignment. The headset provides articulated hand tracking (skeletal joint transforms). Our goal is not only plausible robot motion but also touch alignment: when the participant sees the robot end-effector contacting their hand/arm, they should simultaneously feel contact at the corresponding location. We achieve this by overlaying the robot avatar onto the co-located physical interaction space (Section LABEL:subsec:colocation) and solving inverse kinematics (IK) chains to match tracked human hand targets (Figure 2, steps 2–3). Specifically, we use three IK chains: (i) arm-to-palm IK from the Franka arm base to the robot palm pose using the tracked palm as the end-effector target, (ii) thumb fingertip IK from thumb base to thumb tip using the tracked thumb tip pose, and (iii) index fingertip IK from index base to index tip using the tracked index tip pose. This prioritizes precise alignment at the palm and two salient fingertips that frequently initiate and regulate contact, enabling more convincing “robot touch” during fingertip interactions.
Remaining fingers via direct joint-angle mapping. For the remaining fingers (middle, ring, little), we perform direct joint-angle mapping: each robot finger joint follows the corresponding human finger joint rotation. We do not solve IK for all fingertips because the Allegro hand is substantially larger than a human hand, making simultaneous absolute alignment of all five fingertip poses difficult without introducing large errors or unnatural postures. In practice, full-fingertip IK tended to destabilize contact alignment during dexterous motion; therefore, we prioritize thumb and index fingertips.
Contact realism with fingertip covers. To further increase realism during fingertip touch, the operator can wear 3D-printed fingertip covers shaped to match the Allegro fingertip geometry (panel 6 in Figure 1), helping the felt contact better match the seen robot fingertip (e.g., taps and light presses).
3. Conducting user studies
To test and validate the proposed framework, we designed a pilot study manipulating factors that are specifically of interest to our research on socio-physical HRI but that can also showcase how the VR2 system can be used for prototyping and conducting user studies in a convenient balance of experimental controllability and interaction realism.
3.1. Exemplary experimental design
The study is meant to investigate how social retargeting can affect user experience and trust in the context of robot-initiated contact. In this domain, a number of studies have been conducted either with videos In a WOZ experiment we aim to show that the experience of (functional/affective) touch by the robot can be improved by a combination of retargeted non-verbal cues. In particular we hypothesize that such experience can be modulated differently depending on the context and the richness of displayed cues, i.e.:
HP1: the more human-like the social behavior, the more likable the robot. To test this hypothesis, we consider 3 levels of non-verbal expressivity: 1) just head motion; 2) head and eye movements; 3) head motion, eye movements, and face expressions.
HP2: the context and function of the touch can also modulate how the robot is perceived. In this sense, we have two contexts: 1) functional; 2) playful. We expect the robot to be perceived as more competent when the contact is administered with a functional reason, rather than for entertainment.
Participants are informed that they will interact with a robot in VR across three trials, each time the robot has different expressive capabilities. In the functional context, the robot introduces itself as a nurse in charge of taking the temperature with a sensor placed on its finger. It then asks the participant to turn up one hand and touches them on the back of the intended hand. The robot explains that it will draw a shape on the palm to perform the measurement and asks to guess the pattern. If the person is wrong, the robot asks for another guess, otherwise it acknowledges the right answer, states that the participant is in good health, and concludes the interaction. For the playful context, the interaction unfolds similarly, but the robot acts as a companion and asks to turn and touch the hand to play a guessing game. In the different trials, the robot just moves the head with the eyes displayed but fixed in position and size (head condition), moves the head and eyes independently (head and eye condition), or displays head motion, gaze direction, and eye/ear expressions as detailed in Section 2.2. Throughout the study, the wizard will be played by the same operator to minimize differences in stimulus administration across participants. The wizard impersonating the robot will follow a precise script, with consistent movements and face expressions, thoroughly rehearsed, across all conditions. That is, across expressivity conditions, the wizard performs always the same behavior, but only the condition-related cues are retargeted and displayed by the robot in VR. Speech lines are always the same within each context scenario and delivered as explained in Section 2.1. Essentially, the wizard is solely responsible for turn-taking, triggering a new utterance only when the participant is ready and behaving consistently in every trial, independent of context and expressivity levels. The two hypotheses will be tested in a mixed design, with participants experiencing only one context, but all expressivity levels.
3.2. Data Collection
By means of the VR2 framework, we can collect timestamped data streams from both headsets during each interaction. In particular, for the participant, at any time point, we collect which part of the robot body (head, trunk, hands) they are focusing their gaze on (intersection of the gaze vector with a robot collider). For the robot/operator, we collect contact events with the participant by detecting collisions between the robot’s hand and the participant’s hand colliders. We further collect all blendshapes from both headsets and compute action units (AUs) according to the Facial Action Coding System (FACS, (Ekman and Friesen, 1978)). We further collect speech events considering pedal key presses. Pose data from both actors (heads, hands or other joints of interest) can also be acquired. All this data can be used to relate specific experimental events in time (robot speech, robot physical contact) to behavioral reactions from the participants (gaze, face expressions) as proxies of attention, acceptance/trust or other attitudes toward the robot (see Figure 3. In aggregate form, such behavioral measurements can be compared across conditions and correlated with subjective measurements, such as questionnaires administered at the end of each interaction, still providing a more granular picture of the dynamic nature of the investigated constructs.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.References
- Facial action coding system. Environmental Psychology & Nonverbal Behavior. Cited by: §3.2.
- The path towards contact-based physical human–robot interaction. Robotics and Autonomous Systems 182, pp. 104829. External Links: Document, ISSN 09218890 Cited by: §1.
- Merging physical and social interaction for effective human-robot collaboration. In 2018 IEEE-RAS 18th International Conference on Humanoid Robots (Humanoids), pp. 1–9. External Links: Document, ISBN 978-1-5386-7283-9 Cited by: §1.
- Charting user experience in physical human-robot interaction. ACM Transactions on Human-Robot Interaction 13, pp. 1–29. External Links: Document, ISSN 25739522 Cited by: §1.
Appendix A Facial-expression mapping equations
We map a subset of Quest Pro blendshape coefficients and gaze direction to the robot screen-face parameters used in Figure 2 (step 1). Let , , , , , and denote the controllable degrees of freedom. The mapping is:
Here, is the maximum gaze angle used for normalization, and limits each component to .